7P A n a r
The position of a block moving outward along the slot in the platform is r = 0.022 t^4 m/s where r is in m and t is in seconds. The platform angular velocity increases at a constant rate of 0.5 rad/s/s. If the block and platform start from rest at r = 0 (the center), determine the magnitudes of the block (a) velocity and (b) acceleration when the block is at position r = 8 m. 0.022 4 0.5 8
Answer #1

Position of the block is = 0.0227 ms angular acceleration is 2 = 0.5 rad. /?. we calculate it when now = 8 m. 8 = 0.022 ++ 0.where wo = o rad guc. (initial angular speed) W = 0 + 0.5 x 4.366 w = 2.183 vtad /sec. fourmala Here we use the radius & anguof acculeration the block is radial 17. qe = (17.464) 9 = 38.123 m/s² acceleration is magnitude of now a = 9,+ qa? ( 38 a = 4